a basket contains 10 apples, of which 5 are rotten. what is the probability that a person who buys 4 apples will get none that are rotten?
Answer by Fombitz(32388)

In other words, what's the probability of getting 4 good apples in a row from a basket of 5 good and 5 rotten apples.
P(1st good apple)=5/10
P(2nd good apple)=4/9
P(3rd good apple)=3/8
P(4th good apple)=2/7
P(4 good apples)=(5/10)(4/9)(3/8)(2/7)=120/5040=1/42